Propylene oxide is a common liquid compound that vaporizes readily. Determine the molar mass of propylene oxide from the following data:
Sample mass = 0.0780 g Volume of gas = 125 mL
Temperature = 0.0 °C Pressure = 183 mm Hg

Molar mass =??? g/mol
1
Expert's answer
2020-04-29T13:29:43-0400

pV=nRT. n= pV/RT= 183 x 125 / 62400 x 273 = 0.00134 moles.

n=m/M. M= m/n= 0.078/0.00134 = 58.2.